OpenStudy (anonymous):

Help when you can:) Which of the following constants can be added to x^2 - 3x to form a perfect square trinomial? A. 1 1/2 B. 2 1/4 C. 4 1/2

OpenStudy (anonymous):

(ax+b)^2 = (ax)^2 + 2abx + b^2 in our case a = 1 2ab = -3 so 2b = -3 (Cause a=1) so b=-3/2 so b^2 = 9/4 which is 2 1/4.

OpenStudy (anonymous):

Thank you so much! @Coolsector